# ChemMLLM: Chemical Multimodal Large Language Model

### v1

@@ Abstract
Multimodal large language models (MLLMs) have made impressive progress in many applications in recent years. However, chemical MLLMs that can handle cross-modal understanding and generation remain underexplored. To fill this gap, in this paper, we propose ChemMLLM, a unified chemical multimodal large language model for molecule understanding and generation. Also, we design five multimodal tasks across text, molecular SMILES strings, and image, and curate the datasets. We benchmark ChemMLLM against a range of general leading MLLMs and Chemical LLMs on these tasks. Experimental results show that ChemMLLM achieves superior performance across all evaluated tasks. For example, in molecule image optimization task, ChemMLLM outperforms the best baseline (GPT-4o) by 118.9% (4.27 vs 1.95 property improvement). The code is publicly available at https://github.com/bbsbz/ChemMLLM.git.
ChemMLLM: Chemical Multimodal Large Language Model
Qian Tan1,2, Dongzhan Zhou2,†, Peng Xia2, Wanhao Liu1,2, Wanli Ouyang2, Lei Bai2, Yuqiang Li2,†, Tianfan Fu3,2,† 1University of Science and Technology of China 2Shanghai Artificial Intelligence Laboratory 3Nanjing University tanqian@mail.ustc.edu.cn,liyuqiang@pjlab.org.cn,futianfan@gmail.com
@@ 1 Introduction
Figure 1: Overall architecture of ChemMLLM. (a) Image tokenizer and de-tokenizer. The image tokenizer employs CNN to extract spatial feature maps, where each $n_{z}$-dimensional spatial code is quantized into a discrete latent code via vector quantization (VQ). The resulting codebook indices serve as the final image tokens. Image de-tokenizer uses CNN to reconstruct image from discrete feature map. Then, a patch-based discriminator predicts whether the patch is fake (f) or real (r) (Section 3.1); (b) SMILES tokenizer and de-tokenizer. SMILES tokenization is consistent with text, and is mapped into a token sequence via text tokenizer; (c) ChemMLLM Training; (d) ChemMLLM Inference; (e) two-stage training paradigm for ChemMLLM (Section 3.3).
Multimodal large language models (MLLMs) have shown strong abilities in understanding and generating content across text, images, and audio OpenAI (2024); Sun et al. (2024); Team (2024); Liu et al. (2024); Zhou et al. (2024); Xie et al. (2024); Wang et al. (2024), enabling more natural human–AI interaction. Chemistry is inherently multimodal, involving textual descriptions, structured formats like SMILES Weininger (1988) [^1] A SMILES (Simplified Molecular Input Line Entry System) string is a compact, text-based representation of a molecule’s structure that encodes its atomic composition and connectivity in a linear format. , and molecular images. Recent works have demonstrated initial success in adapting MLLMs for chemical applications such as property prediction and reaction analysis Cao et al. (2023); Zhang et al. (2024b); Luo et al. (2024); Li et al. (2025). However, these models largely focus on understanding tasks and treat images only as input. In contrast, molecular visuals are central to how chemists communicate and reason. Enabling image generation from chemical language or structure would greatly expand the expressiveness of MLLMs in chemistry Kosenkov and Kosenkov (2021). Yet, an integrated Chemical MLLM that supports both multimodal understanding and generation for chemistry remains lacking.
The challenges to build such a model are as follows: (1) Vision-based chemistry tasks and datasets remain underexplored. (2) Specificity of molecule image. Unlike natural images, molecule images are sparse, containing large areas of empty space, and are composed strictly of straight lines. General MLLMs result in unclear and distorted molecules (Figure 15). (3) Challenge for selecting an effective framework for chemical MLLM to seamlessly fuse diverse modalities, including discrete SMILES string/text, and continuous molecule images.
@@
Contribution. For ease of exposition, we summarize our main contribution as:
  • A chemical multimodal LLM understanding and generating molecule in a unified framework. We propose and implement ChemMLLM, the first unified model to understand and generate molecules in text, SMILES, and image modality to the best of our knowledge (Section 3).
  • A multimodal chemical dataset suite. We develop five new datasets to train and evaluate the chemical multimodal capability of MLLMs, which encompass a diverse spectrum of multimodal processing (Section 4).
  • A variety of reliable evaluation. We benchmark the performance of different models on our proposed five tasks and ChemMLLM achieves dominating performance. For example, in molecule image optimization (image-to-image) tasks, ChemMLLM outperforms the best baseline (GPT-4o) by 118.9%, achieving a logP (optimizing property) increase of 4.27 compared to 1.95 (best baseline, GPT-4o) (Section 5).
@@ 2 Related Work
Qwen-VL-Chat 0.09 ±$\pm 0.001$ 0.01 ±$\pm 0.0009$ 0.32 ±$\pm 0.003$ 0.07 ±$\pm 0.001$ 0.22 ±$\pm 0.002$ 0.19 ±$\pm 0.001$
InternVL-Chat-v1.5 0.04±$\pm 0.0008$ 0.001 ±$\pm 0.0002$ 0.38±$\pm 0.003$ 0.07 ±$\pm 0.001$ 0.26 ±$\pm 0.002$ 0.19±$\pm 0.001$
LLaVA-v1.5-7B 0.08 ±$\pm 0.001$ 0.004 ±$\pm 0.0004$ 0.33 ±$\pm 0.002$ 0.07 ±$\pm 0.001$ 0.23 ±$\pm 0.002$ 0.20 ±$\pm 0.001$
GPT-4o 0.16 ±$\pm 0.003$ 0.07±$\pm 0.002$ 0.28±$\pm 0.005$ 0.13±$\pm 0.003$ 0.23±$\pm 0.004$ 0.22±$\pm 0.004$
ChemVLM-8B 0.15 ±$\pm 0.004$ 0.08 ±$\pm 0.003$ 0.28±$\pm 0.006$ 0.13 ±$\pm 0.003$ 0.23 ±$\pm 0.005$ 0.23 ±$\pm 0.005$
ChemMLLM (ours) 0.33 ±$\pm 0.005$* 0.21±$\pm 0.005$* $\pm 0.005$* 0.31±$\pm 0.006$* 0.43±$\pm 0.005$* 0.43±$\pm 0.005$*
Multimodal LLM (MLLM)
With the advancement of large language models and multimodal learning, numerous high-performing multimodal large language models (MLLMs) have recently emerged. Notable MLLMs include GPT-4o OpenAI (2024), which extends GPT-4V OpenAI (2023) to understand and generate across different modalities, including text, image, and audio. Emu-3 Wang et al. (2024) unifies vision understanding and generation via discrete token modeling. Chameleon Team (2024) aligns modalities at the token level for flexible multimodal generation, while LlamaGen Sun et al. (2024) treats images as language-like sequences for scalable autoregressive image generation. Lumina-mGPT Liu et al. (2024) trains a family of models to generate flexible photorealistic images from text descriptions based on Chameleon. Transfusion Zhou et al. (2024) and Show-o Xie et al. (2024) combine diffusion and transformer for multimodal understanding and generation. For vision understanding, well-known models include LLaVA Liu et al. (2023), BLIP-2 Li et al. (2023), Qwen-VL-Chat Bai et al. (2023), InternVL-Chat Chen et al. (2023), MiniGPT-4 Zhu et al. (2023), Gemini Team et al. (2023), Flamingo Alayrac et al. (2022) and OpenFlamingo Awadalla et al. (2023). Although current MLLMs perform well across modalities, these models struggle with chemical tasks due to a misalignment between general and domain-specific knowledge.
Chemical LLM
MLLMs have also exhibited strong potential in addressing chemistry-related tasks, particularly in bridging the modality gap between textual descriptions and molecular representations. Concretely, Instruct-Mol Cao et al. (2023) and MV-Mol Luo et al. (2024) utilize LLaVA’s architecture Liu et al. (2023) and Q-former Li et al. (2023) to align molecular structure and text modality, respectively. ChemLLM Zhang et al. (2024a) uses a high-quality chemical dataset to fine-tune InternLM2 Cai et al. (2024). ChemVLM Li et al. (2025) extends ChemLLM Zhang et al. (2024a) to understand images by adopting a projector-based method to align vision information and text information. UniMoT Zhang et al. (2024b) uses a molecule tokenizer to align the graph modality molecule with text. Despite progress in chemical tasks, existing models can not generate molecular images, limiting their utility in more intuitive, visual forms of interaction. To fill this gap, We propose ChemMLLM, a unified framework that understands and generates molecules in text, SMILES, and image formats.
3 Method
@@
Qwen-VL-Chat 0.78 ±$\pm 0.02$ 7073.3 ±$\pm 858.2$ 58.0 ±$\pm 1.8$ 0.14 ±$\pm 0.02$ 4.4 ±$\pm 0.46$ 1.4 ±$\pm 0.04$ 0.19 ±$\pm 0.02$ $>$1e+4 ±$\pm 563.6$ 82.8 ±$\pm 1.9$ 99.3%
InternVL-Chat-v1.5 0.59 ±$\pm 0.02$ $>$1e+4 ±$\pm 1595.5$ 83.6±$\pm 3.6$ 0.04±$\pm 0.05$ 9.2±$\pm 0.62$ 2.3±$\pm 0.08$ 0.29±$\pm 0.03$ 2158.8±$\pm 432.6$ 28.7 ±$\pm 1.5$ 55.0%
[…]
ChemVLM-8B 0.84 ±$\pm 0.02$ 9573.4 ±$\pm 1992.9$ 56.9 ±$\pm 4.5$ 0.38 ±$\pm 0.05$ 4.9 ±$\pm 0.58$ 1.6 ±$\pm 0.08$ 0.26 ±$\pm 0.06$ 5332.0 ±$\pm 747.5$ 53.66 ±$\pm 2.7$ 31.3%
ChemMLLM (ours) 0.97±$\pm 0.004$* 789.7±$\pm 162.2$* 16.1±$\pm 0.72$* 0.92±$\pm 0.01$* 0.70±$\pm 0.14$* 0.52±$\pm 0.02$* 0.97±$\pm 0.005$* 152.6±$\pm 39.2$* 6.0±$\pm 0.33$* 99.6%
Hbd Hba Rb QED
Qwen-VL-Chat -0.02 ±$\pm 0.008$ 4.8 ±$\pm 0.75$ 1.4±$\pm 0.05$ 0.05±$\pm 0.01$ 30.5±$\pm 1.2$ 4.9±$\pm 0.08$ $\pm 0.1$ 145.4±$\pm 30.8$ 6.8±$\pm 0.31$ $\pm 0.0$
InternVL-Chat-v1.5 0.03±$\pm 0.05$ 9.9±$\pm 0.89$ 2.2±$\pm 0.09$ 0.22±$\pm 0.04$ 10.2±$\pm 0.85$ 2.4±$\pm 0.08$ $\pm 0.04$ $\pm 4.3$ 4.8±$\pm 0.19$ 0.003±$\pm 0.03$ 4.0±$\pm 3.2$ 0.4±$\pm 0.08$
LLaVA-v1.5-7B 0.004 ±$\pm 0.05$ 53.3±$\pm 27.3$ 3.7±$\pm 0.33$ 0.04±$\pm 0.05$ 23.7±$\pm 5.4$ 2.9±$\pm 0.19$ 0.03 ±$\pm 0.04$ 39.9±$\pm 10.4$ 3.8±$\pm 0.26$0.11±$\pm 0.08$ >>$\pm>1e+5$ 24.5±$\pm 18.6$
ChemVLM-8B 0.49 ±$\pm 0.09$ 4.2 ±$\pm 0.85$ 1.3 ±$\pm 0.08$ 0.32 ±$\pm 0.07$ 27.2 ±$\pm 1.95$ 4.5 ±$\pm 0.14$ 0.10 ±$\pm 0.06$ 45.8 ±$\pm 5.02$ 5.5 ±$\pm 0.22$0.003 ±$\pm 0.06$ 0.24 ±$\pm 0.02$ 0.37 ±$\pm 0.01$
ChemMLLM (ours) 0.96±$\pm 0.004$* 0.18±$\pm 0.02$* 0.13±$\pm 0.01$* 0.94±$\pm 0.007$* 0.79±$\pm 0.12$* 0.44±$\pm 0.02$* 0.94±$\pm 0.01$* 1.6±$\pm 0.33$* 0.59±$\pm 0.03$* 0.91±$\pm 0.006$* 0.008±$\pm 0.0004$* 0.06±$\pm 0.002$*
Table 3: Results on img2property task: Hbd, Hba, Rb, and QED (best, 2nd best, *: significantly better (statistically)).
Overview. Our framework uses an image tokenizer to transfer images into discrete tokens, aligning with texts and molecule’s SMILES strings at the token level, known as “Image Tokenizer-LLM-Image De-tokenizer” architecture Team (2024). First, Section 3.1 discusses the molecule image tokenizer. Then, Section 3.2 describes how to combine images, texts and SMILES strings. Finally, the training strategy is discussed in Section 3.3. The whole pipeline is shown in Figure 1. For ease of understanding, we list key mathematical notations in Table 10 in Appendix.
3.1 Mol-VQGAN for Molecule Image Generation
Following Chameleon Team (2024), to enable multimodal information alignment in the early stage, images need to be discretized into token sequences similar to text. To proceed it, we train Vector Quantized Generative Adversarial Network (VQGAN) Esser et al. (2021) on molecule images, known as Mol-VQGAN. Mol-VQGAN compresses images into a discrete latent space, which aligns well with the sequential nature of text modeling in large language models. Specifically, Mol-VQGAN uses Vector Quantized Variational Auto-Encoder (VQVAE) Van Den Oord et al. (2017) as the generator and patch-based discriminator Isola et al. (2017). VQVAE compresses images into discrete spaces and reconstructs images from that space. Mol-VQGAN adds a discriminator and perceptual loss to VQVAE to keep good perceptual quality and improve the performance. Formally, Mol-VQGAN takes an image $x\in\mathbb{R}^{H\times W\times 3}$ ($H/W$ are the height/width of the input image), and transfers it into discrete representations $z_{\textbf{q}}$ by encoding $\hat{z}\in\mathbb{R}^{h\times w\times n_{z}}=E(x)$ ($h/w/n_{z}$ are the height/width/channels of the feature map), and finding the closest codebook entry for each spatial code $\hat{z}_{ij}\in\mathbb{R}^{n_{z}}$ (also known as vector quantization (VQ) process, denoted $\textbf{q}(\cdot)$):
$\hat{x}=G(z_{q})=G(\textbf{q}(E(x))).$ (2)
The discretization and reconstruction process is illustrated in Figure 1(a).
The VQGAN’s objective function is:
$\min_{E,G,Z}\max_{D}\big{[}\mathcal{L}_{vqvae}(E,G,Z)$ (3)
$+\lambda_{1}\mathcal{L}_{perceptual}(E,G,Z)$
$+\lambda_{2}\mathcal{L}_{GAN}(\{E,G,Z\},D)\big{]},$
where (1) the first term $\mathcal{L}_{vqvae}(E,G,Z)=||x-G(\hat{z}-sg(\hat{z}-z_{q}))||_{2}^{2}+||sg[\hat{z}]-z_{q}||^{2}_{2}+||sg[z_{q}]-\hat{z}||^{2}_{2}$ is VQVAE loss, where $\mathcal{L}_{rec}=||x-G(\hat{z}-sg(\hat{z}-z_{q}))||_{2}^{2}$ is reconstruction loss. Since the quantization process is non-differentiable, a stop-gradient operation $sg[\cdot]$ is used so that the forward pass operates on quantized vectors, whereas the backward pass leverages continuous vectors for gradient computation Van Den Oord et al. (2017); (2) the second term $\mathcal{L}_{perceptual}(E,G,Z)=||P(x)-P(G(\hat{z}-sg(\hat{z}-z_{q})))||_{2}^{2}$ is perceptual loss, $P$ denotes a perceptual model like Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018), which is used to extract the high-level semantic features; (3) the third term $\mathcal{L}_{GAN}(\{E,G,Z\},D)=\log D(x)+\log(1-D(\hat{z}-sg(\hat{z}-z_{q})))$ is GAN loss, $D$ is a patch-based discriminator Isola et al. (2017) aiming to differentiate original and reconstructed images. $\lambda_{1}$ is a hyperparameter and $\lambda_{2}$ is an adaptive weight computed dynamically to balance the weight of $\mathcal{L}_{GAN}$ and stabilize training, following VQGAN Esser et al. (2021).
3.2 ChemMLLM
MW LogP TPSA
InternVL-Chat-v1.5 0.04±$\pm 0.05$ $>1e+5\pm>1e+5$ 210.5±$\pm 29.4$0.02±$\pm 0.11$ 23.4±$\pm 9.3$ 2.96±$\pm 0.31$ 0.04±$\pm 0.11$ $>1e+4\pm 6205.6$ 58.9±$\pm 8.1$ 75.0%
LLaVA-v1.5-7B 0.07±$\pm 0.27$ $>2e+6\pm>9e+5$ 894.9±$\pm 285.0$0.52±$\pm 0.16$ 1104.4±$\pm 522.3$ 18.0±$\pm 6.7$ 0.48 ±$\pm 0.27$ $>1e+5\pm>1e+5$ 146.3±$\pm 88.5$ 8.5%
GPT-4o 0.77±$\pm 0.04$ 7633.2±$\pm 1954.0$ 55.7±$\pm 4.7$* 0.48±$\pm 0.07$ 5.7±$\pm 0.87$ 1.7±$\pm 0.11$ 0.69±$\pm 0.04$ 1209.2±$\pm 211.1$ 24.3±$\pm 1.7$ 99.0%
ChemLLM-7B-Chat -0.25±$\pm 0.08$ $>7e+4\pm>1e+4$ 244.8±$\pm 16.2$ 0.0005±$\pm 0.05$ 10.0±$\pm 1.2$ 2.7±$\pm 0.19$0.35±$\pm 0.09$ 7666.0±$\pm 937.2$ 80.6±$\pm 4.3$ 35.0%
ChemVLM-8B 0.14±$\pm 0.15$ $>2e+5\pm>1e+5$ 197.39±$\pm 30.9$ $\pm 0.08$ 9.0±$\pm 1.5$ 2.2±$\pm 0.14$ $\pm 0.22$ $>2e+4\pm>2e+4$ 62.5±$\pm 11.0$ 100.0%
ChemMLLM (ours) 0.71±$\pm 0.05$ $>\underline{3e+4}\pm>1e+4$ 119.5±$\pm 12.1$ 0.42±$\pm 0.05$ 13.2±$\pm 9.0$ 1.8±$\pm 0.26$ 0.71±$\pm 0.05$ 1191.6±$\pm 183.6$ 26.5±$\pm 1.91$ 65.5%
@@ 3.2 ChemMLLM
InternVL-Chat-v1.5 -0.1±$\pm 0.08$ 7.3±$\pm 3.8$ 1.5±$\pm 0.18$ 0.09±$\pm 0.08$ 55.3±$\pm 31.1$ 3.4±$\pm 0.52$ 0.06±$\pm 0.08$ 44.9±$\pm 15.5$ 4.3±$\pm 0.41$ 0.07±$\pm 0.08$ $\pm 0.009$ 0.24±$\pm 0.01$
LLaVA-v1.5-7B $\pm 0.0$ 0.44±$\pm 0.27$ 529.7±$\pm 495.4$ 9.3±$\pm 5.0$0.24±$\pm 0.26$ 5789.7±$\pm 3515.2$ 37.5±$\pm 15.7$ 0.11±$\pm 0.24$ 0.10±$\pm 0.03$ 0.25±$\pm 0.04$
GPT-4o 0.58±$\pm 0.07$ 1.8±$\pm 0.35$ 0.86±$\pm 0.07$ 0.57±$\pm 0.07$ 6.2±$\pm 1.0$ 1.7±$\pm 0.12$ 0.45±$\pm 0.07$ 11.9±$\pm 2.1$ 2.4±$\pm 0.17$ 0.59±$\pm 0.04$* $\pm 0.003$ 0.15±$\pm 0.008$
ChemLLM-7B-Chat -0.23±$\pm 0.07$ 3.4±$\pm 0.82$ 1.5±$\pm 0.14$0.22±$\pm 0.06$ 30.3±$\pm 4.3$ 4.9±$\pm 0.31$0.22±$\pm 0.07$ 46.9±$\pm 6.5$ 5.9±$\pm 0.43$0.07±$\pm 0.03$ 0.07±$\pm 0.009$ 0.24±$\pm 0.01$
ChemVLM-8B 0.47±$\pm 0.30$ 6.0±$\pm 0.99$ 1.7±$\pm 0.12$ 0.02±$\pm 0.13$ 75.6±$\pm 58.4$ 3.49±$\pm 0.56$ 0.44±$\pm 0.29$ 25.8±$\pm 4.3$ 3.7±$\pm 0.24$ 0.10±$\pm 0.06$ 0.08±$\pm 0.007$ 0.24±$\pm 0.01$
ChemMLLM (ours) 0.45±$\pm 0.08$ 1.5±$\pm 0.32$ 0.83±$\pm 0.08$ 0.66±$\pm 0.06$ 5.8±$\pm 0.84$ 1.8±$\pm 0.13$ 0.62±$\pm 0.05$* 17.3±$\pm 8.1$ 2.4±$\pm 0.29$ 0.34±$\pm 0.08$ 0.08±$\pm 0.009$ 0.24±$\pm 0.01$

### v2

@@ Abstract
Multimodal large language models (MLLMs) have made impressive progress in many applications in recent years. However, chemical MLLMs that can handle cross-modal understanding and generation remain underexplored. To fill this gap, we propose ChemMLLM, a unified chemical multimodal large language model for molecule understanding and generation. Also, we design five multimodal tasks across text, molecular SMILES strings, and image, and curate the datasets. We benchmark ChemMLLM against a range of general leading MLLMs and Chemical LLMs on these tasks. Experimental results show that ChemMLLM achieves superior performance across all evaluated tasks. For example, in molecule image optimization task, ChemMLLM outperforms the best baseline (GPT-4o) by 116.75% (4.27 vs 1.97 property improvement). The code is publicly available at https://anonymous.4open.science/r/ChemMLLM-0D98/.
@@ 1 Introduction
Recently, multimodal large language models (MLLMs) have demonstrated strong capabilities in understanding and generating across modalities such as text, images, and audio (OpenAI 2024; Sun et al. 2024; Team 2024; Liu et al. 2024; Zhou et al. 2024; Xie et al. 2024; Wang et al. 2024), enabling more natural and intuitive human–AI interaction. Chemistry is inherently multimodal, encompassing textual descriptions, structured formats like SMILES (Weininger 1988) [^1] A SMILES (Simplified Molecular Input Line Entry System) string is a compact, text-based representation of a molecule’s structure that encodes its atomic composition and connectivity in a linear format., and molecular images. Traditional machine learning models have exhibited success in specific chemical tasks (Goh et al. 2017; Fu, Xiao, and Sun 2020), but they are inherently task-specific and lack the capacity for interactive or multimodal reasoning. For example, in the molecule optimization task (the goal is to generate a similar molecule with more desirable properties), classical sequence-to-sequence models (Cho et al. 2014) may generate a SMILES string with improved properties, but cannot produce interpretable visualizations or generalize beyond that task (Figure 1). Recent efforts have begun to adapt MLLMs to chemical applications, such as property prediction and reaction understanding (Cao et al. 2023; Zhang et al. 2024b; Luo et al. 2024; Li et al. 2025). However, these models primarily treat images as inputs and focus on understanding tasks, lacking the ability to generate chemical visuals—an essential part of how chemists communicate and interpret molecular structures (Kosenkov and Kosenkov 2021). An integrated Chemical MLLM that supports both multimodal understanding and generation for chemistry remains lacking.
Figure 1: Motivation. Comparison between task-specific sequence-to-sequence (Seq2Seq) (Cho et al. 2014) model and a unified chemical large language model.
Figure 2: Overall architecture of ChemMLLM. (a) Image&SMILES tokenizer and de-tokenizer. The image tokenizer employs CNN to extract spatial feature maps, where each $n_{z}$-dimensional spatial code is quantized into a discrete latent code via vector quantization (VQ). The resulting codebook indices serve as the final image tokens. Image de-tokenizer uses CNN to reconstruct image from discrete feature map. Then, a patch-based discriminator predicts whether the patch is fake (f) or real (r); SMILES tokenization is consistent with text, which is mapped into a token sequence via text tokenizer; (b) ChemMLLM training; (c) ChemMLLM inference; (d) two-stage training paradigm.
The challenges in building such a model are as follows: (1) Vision-based chemistry tasks and datasets remain underexplored. (2) Specificity of molecule image. Unlike natural images, molecule images are sparse, containing large areas of empty space, and are composed strictly of straight lines. General MLLMs result in unclear and distorted molecules (Figure 17). (3) Challenge for selecting an effective framework for chemical MLLM to seamlessly fuse diverse modalities, including discrete SMILES string/text, and continuous molecule images.
@@
Contribution. Our main contributions include:
  • A chemical multimodal LLM understanding and generating molecules in a unified framework. We propose and implement ChemMLLM, the first unified model to understand and generate molecules in text, SMILES, and image modality, to the best of our knowledge.
  • A multimodal chemical dataset suite. We develop five new datasets to train and evaluate the multimodal chemical capability of MLLMs, which encompass a diverse spectrum of multimodal processing.
  • Thorough empirical studies. We benchmark the performance of different models on our proposed five tasks and ChemMLLM achieves dominating performance. For example, in molecule image optimization (image-to-image) tasks, ChemMLLM outperforms the best baseline (GPT-4o) by 116.75%, achieving a logP (optimizing property) increase of 4.27 compared to 1.97 (best baseline, GPT-4o).
@@ 2 Related Work
InternVL-Chat-v1.5 (20B) 0.04±0.05$\pm$ 0.05± 0.05 $>1e+5\pm>1e+5$ 210.52±30.14$\pm$ 30.14± 30.14 -0.02±0.11$\pm$ 0.11± 0.11 23.46±9.25$\pm$ 9.25± 9.25 2.96±0.31$\pm$ 0.31± 0.31 0.04±0.11$\pm$ 0.11± 0.11 $>1e+4\pm 6269.06$ 58.92±8.26$\pm$ 8.26± 8.26 75.0%
LLaVA-v1.5-7B 0.07±0.27$\pm$ 0.27± 0.27 $>2e+6\pm>9e+5$ 894.97±278.19$\pm$ 278.19± 278.19 -0.52±0.16$\pm$ 0.16± 0.16 1104.41±512.59$\pm$ 512.59± 512.59 18.08±6.78$\pm$ 6.78± 6.78 0.48 ±0.27$\pm$ 0.27± 0.27 $>1e+5\pm>1e+5$ 146.38±89.08$\pm$ 89.08± 89.08 8.5%
GPT-4o 0.77±0.04$\pm$ 0.04± 0.04 7633.20±1901.91$\pm$ 1901.91± 1901.91 55.74±4.67$\pm$ 4.67± 4.67 0.48±0.08$\pm$ 0.08± 0.08 5.79±0.87$\pm$ 0.87± 0.87 1.79±0.11$\pm$ 0.11± 0.11 0.69±0.04$\pm$ 0.04± 0.04 1209.22±206.85$\pm$ 206.85± 206.85 24.34±1.78$\pm$ 1.78± 1.78 99.0%
ChemLLM-7B-Chat -0.25±0.08$\pm$ 0.08± 0.08 $>7e+4\pm>1e+4$ 244.86±15.96$\pm$ 15.96± 15.96 0.0005±0.05$\pm$ 0.05± 0.05 10.00±1.27$\pm$ 1.27± 1.27 2.78±0.19$\pm$ 0.19± 0.19 -0.35±0.09$\pm$ 0.09± 0.09 7666.05±950.60$\pm$ 950.60± 950.60 80.66±4.35$\pm$ 4.35± 4.35 35.0%
ChemVLM-8B 0.14±0.15$\pm$ 0.15± 0.15 $>2e+5\pm>1e+5$ 197.39±31.37$\pm$ 31.37± 31.37 $\pm 0.08$ 9.01±1.59$\pm$ 1.59± 1.59 2.21±0.14$\pm$ 0.14± 0.14 $\pm 0.22$ $>2e+4\pm>2e+4$ 62.55±11.02$\pm$ 11.02± 11.02 100.0%
ChemMLLM-7B (ours) 0.71±0.05$\pm$ 0.05± 0.05 $>3e+4\pm>1e+4$ $\pm 11.95$ 0.42±0.05$\pm$ 0.05± 0.05 13.26±9.18$\pm$ 9.18± 9.18 1.89±0.26$\pm$ 0.26± 0.26 0.71±0.05$\pm$ 0.05± 0.05 1191.67±178.69$\pm$ 178.69± 178.69 26.57±1.91$\pm$ 1.91± 1.91 65.5%
ChemMLLM-34B (ours) 0.81±0.03$\pm$ 0.03± 0.03 5241.95±>955.60$\pm$>955.60± > 955.60 48.75±4.00$\pm$ 4.00± 4.00 0.54±0.07$\pm$ 0.07± 0.07 3.70±1.24$\pm$ 1.24± 1.24 1.26±0.10$\pm$ 0.10± 0.10* 0.84±0.02$\pm$ 0.02± 0.02* 389.86±46.03$\pm$ 46.03± 46.03* 15.69±0.89$\pm$ 0.89± 0.89* 90.0%
@@
InternVL-Chat-v1.5 (20B) -0.10±0.08$\pm$ 0.08± 0.08 7.32±3.86$\pm$ 3.86± 3.86 1.55±0.18$\pm$ 0.18± 0.18 0.09±0.08$\pm$ 0.08± 0.08 55.30±30.77$\pm$ 30.77± 30.77 3.44±0.52$\pm$ 0.52± 0.52 0.06±0.08$\pm$ 0.08± 0.08 44.96±15.67$\pm$ 15.67± 15.67 4.31±0.41$\pm$ 0.41± 0.41 0.07±0.08$\pm$ 0.08± 0.08 0.09±0.009$\pm$ 0.009± 0.009 0.24±0.01$\pm$ 0.01± 0.01
LLaVA-v1.5-7B $\pm 0.0$ 0.44±0.27$\pm$ 0.27± 0.27 529.70±477.34$\pm$ 477.34± 477.34 9.35±5.16$\pm$ 5.16± 5.16 -0.24±0.26$\pm$ 0.26± 0.26 5789.76±3598.21$\pm$ 3598.21± 3598.21 37.52±15.98$\pm$ 15.98± 15.98 0.11±0.24$\pm$ 0.24± 0.24 0.10±0.03$\pm$ 0.03± 0.03 0.25±0.04$\pm$ 0.04± 0.04
GPT-4o 0.58±0.07$\pm$ 0.07± 0.07 1.83±0.35$\pm$ 0.35± 0.35 0.86±0.07$\pm$ 0.07± 0.07 0.57±0.07$\pm$ 0.07± 0.07 6.20±1.02$\pm$ 1.02± 1.02 1.75±0.12$\pm$ 0.12± 0.12 0.45±0.07$\pm$ 0.07± 0.07 11.94±2.17$\pm$ 2.17± 2.17 2.45±0.17$\pm$ 0.17± 0.17 0.59±0.04$\pm$ 0.04± 0.04 0.03±0.003$\pm$ 0.003± 0.003 0.15±0.008$\pm$ 0.008± 0.008
ChemLLM-7B-Chat -0.23±0.07$\pm$ 0.07± 0.07 3.46±0.82$\pm$ 0.82± 0.82 1.50±0.14$\pm$ 0.14± 0.14 -0.22±0.06$\pm$ 0.06± 0.06 30.31±4.44$\pm$ 4.44± 4.44 4.91±0.31$\pm$ 0.31± 0.31 -0.22±0.07$\pm$ 0.07± 0.07 46.90±6.50$\pm$ 6.50± 6.50 5.96±0.43$\pm$ 0.43± 0.43 -0.07±0.03$\pm$ 0.03± 0.03 0.07±0.009$\pm$ 0.009± 0.009 0.24±0.01$\pm$ 0.01± 0.01
ChemVLM-8B 0.47±0.30$\pm$ 0.30± 0.30 6.05±0.99$\pm$ 0.99± 0.99 1.70±0.12$\pm$ 0.12± 0.12 0.02±0.13$\pm$ 0.13± 0.13 75.68±59.20$\pm$ 59.20± 59.20 3.49±0.56$\pm$ 0.56± 0.56 0.44±0.29$\pm$ 0.29± 0.29 25.89±4.29$\pm$ 4.29± 4.29 3.78±0.24$\pm$ 0.24± 0.24 0.10±0.06$\pm$ 0.06± 0.06 0.08±0.007$\pm$ 0.007± 0.007 0.24±0.01$\pm$ 0.01± 0.01
ChemMLLM-7B (ours) 0.45±0.08$\pm$ 0.08± 0.08 1.58±0.32$\pm$ 0.32± 0.32 0.83±0.08$\pm$ 0.08± 0.08 0.66±0.06$\pm$ 0.06± 0.06 5.87±0.84$\pm$ 0.84± 0.84 1.87±0.13$\pm$ 0.13± 0.13 0.62±0.05$\pm$ 0.05± 0.05* 17.37±8.33$\pm$ 8.33± 8.33 2.48±0.29$\pm$ 0.29± 0.29 0.34±0.08$\pm$ 0.08± 0.08 0.08±0.009$\pm$ 0.009± 0.009 0.24±0.01$\pm$ 0.01± 0.01
ChemMLLM-34B (ours) 0.72±0.05$\pm$ 0.05± 0.05 0.91±0.11$\pm$ 0.11± 0.11* 0.63±0.05$\pm$ 0.05± 0.05* 0.81±0.02$\pm$ 0.02± 0.02* 2.00±0.19$\pm$ 0.19± 0.19* 1.08±0.06$\pm$ 0.06± 0.06* 0.74±0.03$\pm$ 0.03± 0.03* 7.85±2.48$\pm$ 2.48± 2.48 1.85±0.15$\pm$ 0.15± 0.15* 0.56±0.05$\pm$ 0.05± 0.05 0.04±0.005$\pm$ 0.005± 0.005 0.17±0.01$\pm$ 0.01± 0.01